# Retrieve, Merge, Predict: Augmenting Tables with Data Lakes

### v6

@@ 4 Implementations of the retrieve-merge-predict pipeline
We now discuss the pipeline steps in Figure 1. For the Retrieve step, we discuss different possible retrieval strategies; for the Merge step, we propose different join selectors and aggregation methods, and we explain how these two parts are intertwined; finally, for the Predict step, we go over the ML methods used to test the prediction performance. We explore each section in more detail in Appendix B.
@@
Retrieval methods usually involve offline preparation steps such as building sketches (Zhu et al., 2023; Fernandez et al., 2019), indices (Zhu et al., 2019), creating embeddings (Fan et al., 2023b), or profiling tables in the data lake (e.g., measuring the containment); preparation includes persisting the data structures for future use. Afterwards, candidate joins can be obtained by querying the retrieval method.
For retrieval strategies that rely on Jaccard containment (Equation 1), a candidate column is a good join if its containment is large, i.e. the candidate column contains a large fraction of the entities of the query column. Alternative strategies may rely on different metrics, such as cosine similarity between embeddings (Fan et al., 2023b).
We consider four retrieval methods that represent different levels of query cost to better observe the resource expense-effectiveness trade-off. Approximate methods are represented by MinHashLSHEnsemble (MinHash) (Zhu et al., 2016), which builds an index on the data lake, and when queried returns candidates whose Jaccard containment is greater than a certain threshold without ranking them. Conversely, Exact Matching measures the exact Jaccard containment between a query column and every column in the data lake. We propose Hybrid MinHash as a combination of Exact Matching and MinHash that takes the candidates returned by MinHash and measures their exact containment, thus limiting the number of columns to be considered and reducing the computational cost. Finally, Starmie (Fan et al., 2023b) builds embeddings for the query column and every column in the data lake using a language model, then combines cosine and Jaccard similarity to rank candidates. More detail is provided in Appendix subsection B.1.
The rest of the pipeline is transparent to the retrieval method: given a list of join candidates, it will test each candidate regardless of how the list was constructed. To reflect practical limitations on runtime and compute resources, we introduce a budget constraint on the number of retrieved candidates to evaluate. We choose to keep the top-30 candidates, as we observe that almost all candidates with high containment are within this limit (Figure 8 in Appendix), and it has empirical confirmation in practice (Spotify, 2020). In Section 5, we demonstrate how even such a small budget can cause problems with scalability.
@@
We consider two different overarching strategies for selecting candidates: metric-based selectors, which rely on a metric to choose which candidates should be joined, and results-based selectors, which instead iterate over candidates and select the best by explicitly executing joins. The first category includes Highest Containment, which ranks candidates by containment, then joins the first, and Top-$k$ Full Join, which relies entirely on the ranking provided by the retrieval method to join the first $k$ candidates without filtering them. For the second category, we implement Best Single Join and Stepwise Greedy Join. The first works with one candidate at each iteration, during which it first joins the candidate, then it trains and evaluates a prediction model on the resulting table; after iterating over all candidates, it selects the single candidate with the best performance. The second method iterates over each candidate like in the previous case, however it retains all candidates that improve the prediction performance, so that the augmented table grows over time as new candidates are joined. Stepwise Greedy Join is related to forward feature selection (Guyon & Elisseeff, 2003), as it greedily adds new tables as features during each iteration, and represents a common approach in the database literature (Galhotra et al., 2023; Chepurko et al., 2020). Appendix subsection B.2 details the implementations.
@@
When joining tables for a downstream ML task, one-to-many relationships must be aggregated to avoid replicating samples in the base table, which would modify the initial data sampling. We test three aggregation strategies: Any selects any row at random from each group of matched tuples; Mean replaces all duplicated numerical (categorical) values by the mean (most frequent) of all values for that attribute in the group; finally DFS (Deep Feature Synthesis (Kanter & Veeramachaneni, 2015)) greedily generates new features for each column in the augmented table by aggregating groups of tuples, measuring statistics over the groups (e.g., mean, median, most frequent value), and adding said statistics as features; this is done for every join. We expand on this in Appendix subsection B.3.
Supervised learning with a ML model. The learning step is performed on the final integrated table on the training split returned by the join selector and evaluated on the integrated table on the test split.
We evaluate four predictors. We use a linear baseline: Ridge regressor/classifier with the RidgeCV scikit-learn (Pedregosa et al., 2011) implementation and default parameters. CatBoost (Prokhorenkova et al., 2018) is a state-of-the-art GBDT method, optimized to handle categorical variables. We set the number of iterations to 300. ResNet (Gorishniy et al., 2021) is our baseline neural method. RealMLP (Holzmüller et al., 2024) is a NN-based method that incorporates a number of “good defaults” to improve performance over standard NNs. For ResNet and RealMLP we use the implementation and default parameters provided by pytabkit (Holzmüller et al., 2024). We do not perform hyperparameter optimization.
5 Experimental study
5.1 Settings
Data lakes
We use four YADL variants (Binary, Base, 10k, 50k) (as detailed in Section 3) and Open Data US, a data lake employed in the literature (Galhotra et al., 2023; Zhu et al., 2019; 2016); statistics are reported in Table 1. Starmie did not run on YADL 50k (not enough RAM) and Open Data (noise in the data crashed the model), so these data lakes are excluded from figures that involve Starmie. Appendix Figure 8 reports the measured containment for the base tables used in our experiments. [^2] The data lakes are available at https://doi.org/10.5281/zenodo.10600047, while the code to prepare YADL is at https://github.com/rcap107/YADL and the pipeline is at https://github.com/rcap107/retrieve-merge-predict.
@@ 5.1 Settings
We evaluate five tables from sources external to the lakes: Company Employees, US Elections, 2021 US Road Accidents, and Housing Prices. As “internal” tables, derived from lakes, have been used in several previous works for evaluation (Galhotra et al., 2023; Chepurko et al., 2020), we also include one per lake: US County Population (from YADL), and Schools (from Open Data US). Statistics are reported in Appendix subsection C.8 .
Figure 2: Pareto frontier analysis for the pipeline steps The prediction performance is plotted against run time (top) and peak RAM usage (bottom). Each row presents the same results, broken down by retrieval method (left), join selector method (center) and predictor (right). Each dot represents the average prediction performance and resource cost averaged across base tables and data lakes for a specific configuration (e.g., the leftmost dot in the first row is obtaining by using MinHash, Highest Containment Join and CatBoost). Time for offline retrieval preparation, i.e. indexing and querying, is not reported here.
For all datasets, the values of the query columns must be matched with the entities in YADL using semantic annotation solutions (Huynh et al., 2022; Nguyen et al., 2021). In our experiments, we manually performed the match to remove the noise from this task. Also, query columns are chosen based on what a user may reasonably consider as “key” (e.g., the movie title). We release the matched tables along with the data lakes for reproduciblity.
To reflect the experience of a data scientist that needs to construct a meaningful table starting from a data lake, and to highlight the effect of joins on the downstream task, we run experiments on a depleted version of the tables, i.e., the input tables include only the primary key column and the target column. We provide details on the datasets, computational resources and pipeline implementation in Appendix subsection C.2.
@@ 5.2 Results
Our first goal is to pinpoint which steps of the pipeline have the most significant impact on the studied dimensions: optimizing these influential sections can yield the greater benefits. We rely on Pareto frontier plots (Figure 2) to highlight the cost-performance trade-off for the different steps and configurations.
@@
Based on Figure 2, we define a reference configuration that uses Exact Matching, Best Single Join, aggregation Any, and CatBoost. This configuration represents a good trade-off between performance and compute cost, and it can run on all configurations, base tables and data lakes.
@@
Supervised learning with the ML model is the pipeline step with the starkest difference between methods (Figure 2 right, Table 3): in our scenario, CatBoost is both faster and more effective than its non-tree counterparts. Indeed, CatBoost always outperforms other methods in prediction performance (by up to 26%) and run-time ($\times3$ faster than the slowest competitor), although it has a larger memory footprint.
@@
And yet, while Jaccard containment is an effective first metric, it is not fool proof. We observe experimentally that it does not work well where there is a large degree of redundancy because in such cases many candidates may share the same containment: this is what happens with YADL 50k (Table 2, Appendix Figure 8); a larger computational budget may mitigate this issue, but it is not always an option. Furthermore, it fails in cases where the cardinality of a column is very low (e.g., with columns that contain binary values): this is problematic if the user does not know which columns should be used for querying. In other words, high containment means that we can join without adding too many missing values: this does not guarantee that the added features improve downstream utility.
@@
Pareto frontier analysis (Figure 2, left) shows that Starmie is Pareto optimal for some (resource-intensive) configurations: if resources are available for the offline preparation cost, better querying through Starmie can bring benefits. And yet, the –arguably small– improvements in downstream performance brought by Starmie come with a very large cost in both RAM and offline training time (Figure 5, Appendix Table 6, Appendix Table 10). Furthermore, we had to restrict most of our experiments (including Figure 2) as Starmie could only run on 3 out of our 5 data lakes of interest (Binary, YADL Base and YADL 10k). We had to leave out YADL 50k, where Starmie exceeded our available RAM (150GB), and Open Data which crashed Starmie, probably because it was too noisy. On the other hand, Exact matching is Pareto optimal in most cases, and runs on all data lakes we considered.
With respect to MinHash, Hybrid MinHash brings a substantial improvement in terms of downstream performance (Figure 2, Table 3, Figure 15 ) and top-1 retrieval performance (Appendix Figure 18, Appendix Figure 16). It leverages Exact Matching in order to produce an actual ranking of the candidates. This however comes at the cost of a substantial increase in the query time (Figure 5, Appendix Table 10).
Retrieval: Computation trade-offs
Figure 6: Aggregation: Pareto Frontier Analysis. DFS outperforms other methods, but is much slower to run.
@@
Figure 6 reports aggregation results [^5] More detail is available in Appendix, Figure 13 and Figure 14. The systematic expansion of multiple aggregations in DFS leads to many features, which would make both Full Join and Stepwise Greedy Join intractable, so they are not considered here. The generated features overall bring useful information, as DFS outperforms the simpler aggregations in most cases; however, preparing and employing the new features increases noticeably the training time. This is confirmed by Table 3, which highlights both the improvement in performance ( 2.5%) and the major increase of compute cost ( 6x wrt Any).
Overall results and ablation study
@@
Table 3: Ablation study, prediction difference (higher is better) and runtime difference (lower is better) from the reference configuration. Retrieval Method does not include results from Open Data US and YADL 50k.
Retrieval Method Prediction diff. Runtime diff. Join Selector Prediction diff. Runtime diff.
Exact Matching 0.00% 1.00x Full Join 2.09% 0.57x
Starmie -0.06% 0.98x Stepwise Greedy Join 1.61% 2.91x
Hybrid Minhash -3.19% 0.67x Best Single Join 0.00% 1.00x
Minhash -17.48% 0.44x Highest Containment Join -0.10% 0.12x
Aggregation Prediction diff. Runtime diff. Predictor Prediction diff. Runtime diff.
DFS 2.46% 6.08x Catboost 0.00% 1.00x
Mean 0.02% 1.03x RealMLP -21.24% 32.79x
Any 0.00% 1.00x ResNet -24.11% 10.28x
RidgeCV -26.03% 5.98x
Table 2 reports the median prediction performance across all configurations for a pair “Base table - Data Lake”. From this table it is evident that internal tables (Schools, US County Population) represent a much simpler problem to solve compared to tables that are not sampled from the data lake (all other tables). Table 3 reports the results of an ablation study where we compare the performance and runtime of the reference configuration against the other possible methods; we report the median difference in prediction and runtime.  [^6] This is expanded in Appendix subsection C.2 and Figure 15.
@@ 6 Discussion and conclusions
    Simple metric-based retrieval (Exact matching) and candidate selection (Highest containment) produce comparable or better results than more complex methods (Starmie [^7] Note that our results are not in contradiction with the Starmie study, as it was focused on table unionability, i.e. finding more samples rather feature augmentation which is our focus. and Best single join), while being vastly more efficient (Figure 2 left, center, Table 3).
@@
  5. 5.
    Combining stochastic retrieval with exact metrics (Hybrid MinHash) mitigates some of the drawbacks of basic MinHash and scales well with many query columns (Figure 3(a), Figure 5, Figure 5, Table 3).
The topic that we have studied is at the intersection of database and machine learning research. This intersection still has much open research prospect:

### v7

@@ 4 Implementations of the retrieve-merge-predict pipeline
We now discuss the pipeline steps in Figure 1. For the Retrieve step, we discuss different possible retrieval strategies; for the Merge step, we propose different join selectors and aggregation methods; finally, for the Predict step, we go over the ML methods used to test the prediction performance. We explore each section in more detail in Appendix B.
@@
We consider four retrieval strategies that explore different approaches and trade-offs. We first distinguish between“offline” and “online” methods: methods in the first category require some degree of offline data preparation in order to be executed, such as constructing data structures, or training a specific model; methods in the latter category are instead executed directly “online” as part of the “Data usage” section.
Exact Matching is on “online” method, and measures the exact Jaccard containment between it and every column in the data lake. The “offline” category includes MinHashLSHEnsemble (MinHash) (Zhu et al., 2016), Hybrid MinHash (a method we develop), and Starmie (Fan et al., 2023b). MinHash builds an index on the data lake, and when queried returns candidates whose Jaccard containment is greater than a certain threshold without ranking them. Hybrid MinHash combines Exact Matching and MinHash by taking the candidates returned by MinHashing and measuring their exact containment, thus limiting the number of columns to be considered and reducing the computational cost of Exact Matching, whilst introducing a candidate ranking. Finally, Starmie builds embeddings for the query column and every column in the data lake using a language model, then combines cosine and Jaccard similarity to rank candidates. All methods share the same drawback of requiring some degree of updating/retraining as a data lake evolves over time More detail is provided in Appendix B.1.
The “Data usage” part of the pipeline (1) is transparent to the retrieval method: given a list of join candidates, it will test each candidate regardless of how the list was constructed. To reflect practical limitations on runtime and compute resources, we introduce a budget constraint on the number of retrieved candidates to evaluate. We choose to keep the top-30 candidates for most of our experiments, as we observe that almost all candidates with high containment are within this limit (Figure 7 in Appendix), it has empirical confirmation in practice (Spotify, 2020), and our experiments show that results plateau past this threshold (Figure 4). In Section 5, we explore in detail how the value of $k$ affects scalability and results.
@@
We consider two different overarching strategies for selecting candidates: metric-based selectors, which rely on a metric to choose which candidates should be joined, and results-based selectors, which instead iterate over candidates and select the best by explicitly executing joins. The first category includes Highest Containment, which ranks candidates by containment, then joins the first, and Full Join, which relies entirely on the ranking provided by the retrieval method to join the candidates without filtering them. For the second category, we implement Best Single Join and Stepwise Greedy Join. The first works with one candidate at each iteration, during which it first joins the candidate, then it trains and evaluates a prediction model on the resulting table; after iterating over all candidates, it selects the single candidate with the best performance. The second method iterates over each candidate like in the previous case, however it retains all candidates that improve the prediction performance, so that the augmented table grows over time as new candidates are joined. Stepwise Greedy Join is related to forward feature selection (Guyon & Elisseeff, 2003), as it greedily adds new tables as features during each iteration, and represents a common approach in the database literature (Galhotra et al., 2023; Chepurko et al., 2020). Appendix B.2 details the implementations.
@@
We test three aggregation strategies: Any selects any row at random from each group of matched tuples; Mean replaces all duplicated numerical (categorical) values by the mean (most frequent) of all values for that attribute in the group; finally DFS (Deep Feature Synthesis (Kanter & Veeramachaneni, 2015)) greedily generates new features for each column in the augmented table by aggregating groups of tuples, measuring statistics over the groups (e.g., mean, median, most frequent value), and adding said statistics as features; this is done for every join. We expand on this in Appendix B.3.
Supervised learning with a ML model. The learning step is performed on the final, integrated table obtained by merging the candidates on the training split of the base table. We evaluate four predictors. We use a Ridge regressor/classifier as a linear baseline using the scikit-learn RidgeCV (Pedregosa et al., 2011) implementation and default parameters. CatBoost (Prokhorenkova et al., 2018) is a state-of-the-art GBDT method, optimized to handle categorical variables. ResNet (Gorishniy et al., 2021) is our baseline neural method. RealMLP (Holzmüller et al., 2024) is a NN-based method that incorporates a number of “good defaults” to improve performance over standard NNs. For ResNet and RealMLP we use the implementation and default parameters provided by pytabkit (Holzmüller et al., 2024). We do not perform hyperparameter optimization. Further detail on implementation details, pre-processing of the data, and cross-validation setup are provided in Appendix C.2.
@@
5 Experimental study
5.1 Settings
Data lakes
We use four YADL variants (Binary, Base, 10k, 50k) (as detailed in Section 3) and Open Data US, a data lake employed in the literature (Galhotra et al., 2023; Zhu et al., 2019; 2016); statistics are reported in Table 1. Starmie did not run on YADL 50k (not enough RAM) and Open Data (noise in the data crashed the model), so these data lakes are excluded from figures that involve Starmie. Appendix Figure 7 reports the measured containment for the base tables used in our experiments. [^3] The data lakes are available at https://doi.org/10.5281/zenodo.10600047, the code to prepare YADL is at https://github.com/soda-inria/YADL and the pipeline is at https://github.com/soda-inria/retrieve-merge-predict.
@@ 5.1 Settings
We evaluate four tables from sources external to the lakes: Company Employees, US Elections, 2021 US Road Accidents, and Housing Prices. As “internal” tables, derived from lakes, have been used in several previous works for evaluation (Galhotra et al., 2023; Chepurko et al., 2020), we also include one per lake: US County Population (from YADL) and Schools (from Open Data US). Statistics are reported in Appendix subsection C.7 .
Figure 2: Pareto diagram for the pipeline steps The prediction performance ($y$-axis) is plotted against retrieval + run time (top) and peak RAM usage (bottom). Each row presents the same results, broken down by retrieval method (left), join selector method (center) and predictor (right). Each dot represents the average prediction performance and resource cost averaged across base tables and data lakes for a specific configuration (e.g., the leftmost dot in the first row is obtaining by using MinHash, Highest Containment Join and CatBoost). Time for offline retrieval preparation for MinHash and Starmie is not reported here.
For all datasets, the values of the query columns must be matched with the entities in YADL using semantic annotation solutions (Huynh et al., 2022; Nguyen et al., 2021). In our experiments, we manually performed the match to remove the noise from this task. Also, query columns are chosen based on what a user may reasonably consider as “key” (e.g., the company name). We release the matched tables along with the data lakes for reproduciblity.
To reflect the experience of a data scientist that needs to construct a meaningful table starting from a data lake, and to highlight the effect of joins on the downstream task, we run experiments on a depleted version of the tables, i.e., the input tables include only the primary key column and the target column. We provide details on the datasets, computational resources, cross-validation setup, and pre-processing in Appendix C.2.
@@ 5.2 Results
Table 2: Ablation study, median of the prediction difference (higher is better) and runtime difference (lower is better) from the reference configuration. Retrieval Method does not include results from Open Data US and YADL 50k.
Retrieval Diff. Prediction Diff. Time Selector Diff. Prediction Diff. Time
Exact Matching 0.00 % 1.00x Full Join 2.09 % 0.69x
Starmie -0.06 % 2.20x Stepwise Greedy Join 1.61 % 2.10x
Hybrid MinHash -3.19 % 0.67x Best Single Join 0.00 % 1.00x
MinHash -17.48 % 0.34x Highest Containment Join -0.10 % 0.45x
Aggregation Diff. Prediction Diff. Time Predictor Diff. Prediction Diff. Time
DFS 2.46 % 4.42x CatBoost 0.00 % 1.00x
Mean 0.02 % 1.02x RealMLP -21.24 % 17.52x
Any 0.00 % 1.00x ResNet -24.11 % 5.92x
RidgeCV -26.03 % 3.31x
Our first goal is to pinpoint which steps of the pipeline have the most significant impact on the studied dimensions: optimizing these influential sections can yield the greater benefits. We rely on Pareto optimality plots (Figure 2) to highlight the cost-performance trade-off for the different steps and configurations.
@@
Based on Figure 2, we define a reference configuration that uses Exact Matching, Best Single Join, aggregation Any, and CatBoost. This configuration represents a good trade-off between performance and compute cost, and it can run on all configurations, base tables and data lakes. Table 2 reports an ablation study comparing the performance and runtime of the reference configuration against the other possible methods; we report the median difference in prediction and runtime.  [^4] This is expanded in Appendix subsection C.2 and Figure 14.
@@
Supervised learning with the ML model is the pipeline step with the starkest difference between methods (Figure 2 right, Table 2): in our scenario, CatBoost is both faster and more effective than its non-tree counterparts. Indeed, CatBoost always outperforms other methods in prediction performance (by up to 26%) and run-time (17×17$\times$17 × faster than the slowest competitor).
@@
And yet, while Jaccard containment is an effective first metric, it is not fool proof. We observe experimentally that it does not work well where there is a large degree of redundancy because in such cases many candidates may share the same containment: this is what happens with YADL 50k (Table 3, Appendix Figure 7); a larger computational budget may mitigate this issue, but it is not always an option. In other words, high containment means that we can join without adding too many missing values: this does not guarantee that the added features improve downstream utility.
@@
Pareto diagram (Figure 2, left) shows that Starmie is Pareto optimal for some (resource-intensive) configurations: if resources are available for the offline preparation cost, better querying through Starmie can bring benefits. And yet, the –arguably small– improvements in downstream performance brought by Starmie come with a very large cost in both RAM and offline training time (Figure 2, Figure 22, Appendix Table 6, Appendix Table 10): Starmie has about 7.5×7.5$\times$7.5 × the RAM footprint of the most expensive alternatives; furthermore, it could only run on 3 out of our 5 data lakes of interest (Binary, YADL Base and YADL 10k). On the other hand, Exact matching is Pareto optimal in most cases, and runs on all data lakes we considered.
With respect to MinHash, Hybrid MinHash brings a substantial improvement in terms of downstream performance (Figure 2, Table 2, Figure 14). This comes at the cost of an increase in the query time (Figure 18, Appendix 10), which is however amortized due to the cost of performing cross-validation.
The cost of retrieving candidates by testing many query columns is another important factor in the scalability of the retrieval methods. Indeed, if the user does not know which is the best query column to use, or if there are multiple, they may favor methods that have a shorter query time (such as MinHash and Hybrid MinHash) in order to run the evaluation pipeline faster. We explore scalability in more detail in Appendix C.1.
Figure 4: Trade-off in prediction performance as the number of candidates increases for the Full Join selector. Results are averaged over all base tables and data lakes and represented as a Pareto plot.
Retrieval: performance plateaus quickly as $k$ increases
To understand the impact of $k$ on the prediction performance, we record the prediction performance for a range of values, and report the results in Figure 4. While there is some variability due to the data lake and base table, we observe that results plateau quickly as $k$ increases, saturating for $k$ as low as 8. We explore this facet in more detail in Appendix C.10.
Figure 5: Aggregation: Pareto diagram. DFS outperforms other methods, but is much slower.
@@
Figure 5 reports aggregation results [^5] More detail is available in Appendix, Figure 15. The systematic expansion of multiple aggregations in DFS leads to many features, which would make both Full Join and Stepwise Greedy Join intractable, so they are not considered here. The generated features overall bring useful information, as DFS outperforms the simpler aggregations in most cases; however, preparing and employing the new features increases noticeably the training time. This is confirmed by Table 2, which highlights both the improvement in performance ( 2.5%) and the major increase of compute cost ( 4.42x wrt Any).
Overall results across tables
@@
Table 3 reports the median prediction performance across all configurations for a pair “Base table - Data Lake”. From this table it is evident that internal tables (Schools, US County Population) represent a much simpler problem to solve compared to tables that are not sampled from the data lake (all other tables).
@@ 6 Discussion and conclusions
    Simple metric-based retrieval (Exact matching) and candidate selection (Highest containment) produce comparable or better results than more complex methods (Starmie [^6] Note that our results are not in contradiction with the Starmie study, as it was focused on table unionability, i.e. finding more samples rather than feature augmentation which is our focus. and Best single join), while being vastly more efficient (Figure 2 left, center, Table 2).
@@
    Performance plateaus quickly as the number of retrieved candidates increases, while the resource cost increases steadily (Figure 4).
  5. 5.
@@
  6. 6.
    Combining stochastic retrieval with exact metrics (Hybrid MinHash) mitigates some of the drawbacks of basic MinHash and scales well with many query columns (Figure 3, Table 2).
The topic that we have studied is at the intersection of database and machine learning research. This intersection still has many open research prospect:
@@
Acknowledgements
We acknowledge the invaluable insight and expertise provided by Léo Dreyfus-Schmidt and Du Phan (formerly from Dataiku), without whom the foundations of this work would not have existed. As YADL is based on YAGO 3, we acknowledge the work made by its authors to prepare the original knowledge base, as well as their efforts manually evaluating it. This research was supported by the Project P16, ANR-23-RDIA-0001. We express our gratitude for the financial support that made this study possible. Finally, we would like to thank the anonymous reviewers for their constructive comments and valuable suggestions, which have significantly improved the quality of this manuscript.